### v1

# Structured dataset of reported cloud seeding activities in the United States (2000–2025) using a large language model

Abstract
We present a structured dataset of reported cloud seeding activities in the U.S. from 2000 to 2025, including the year, season, state, seeding agent, apparatus used for deployment, and purpose. Using OpenAI’s o4-mini large language model (LLM), combined with multi-stage PDF-to-text conversion and response-parsing code, we processed 836 historical reports from the National Oceanic and Atmospheric Administration (NOAA) to extract the data. The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo. Our results address the gap in accessible cloud seeding data in the United States, and demonstrate the potential of LLMs to extract structured environmental data from historical documents. More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
Background & Summary
Cloud seeding is a weather modification technique used to enhance precipitation, typically in regions affected by drought, water scarcity, or to support snowpack accumulation.
In the United States, cloud seeding has been practiced since the 1940s but reporting was not standardized until the Weather Modification Reporting Act of 1972 [1]. The Act requires individuals and organizations to notify the U.S. Department of Commerce at least 10 days before and after conducting weather modification operations. These reports are submitted to NOAA via email and compiled on the NOAA Weather Modification website [2].
Currently, NOAA’s publicly available records on their website are available for 2000-2025. We expected earlier records to be available and submitted a data request to NOAA for documents covering 1972–1999, but have not received a response to date. As a result, the dataset currently covers 2000-2025. If older cloud seeding reports from 1972–1999 are recovered and made publicly available, the same pipeline can be applied with minimal modifications to expand the dataset’s coverage.
While the NOAA reports are made publicly available online, they are stored individually as scanned PDFs with inconsistent formatting and structure. This lack of standardization makes it difficult to extract key information at scale, or integrate the reports into other structured datasets. As a result, data on cloud seeding in the U.S., such as locations, dates, purposes, cloud seeding agent, and agent deployment methods, remains largely inaccessible for analysis.
To solve this data access problem, we created a structured dataset from 836 NOAA cloud seeding reports spanning from 2020 to 2025. Our work transforms these scattered, inconsistently formatted documents into a single, structured CSV file that enables the analysis of U.S. weather modification activities over time. The dataset has 94.72% accuracy on all extracted fields when compared with the original PDFs.
This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management—many of which exist in PDF form with highly variable formatting. By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest.
Beyond its methodological contributions, the resulting dataset has potential value across a wide range of applications. Researchers can use it to study long-term patterns in weather modification practices, analyze the use and evolution of different seeding agents and deployment methods, or assess geographic and seasonal trends in operational activity.
For example, while there have been numerous recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g. the 2017 SNOWIE campaign in Western Idaho[3]), there has been little qualitative analysis of reported historical operational cloud seeding activities. With recent renewed interest in both traditional weather modification technology (e.g. the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geo-engineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change, this dataset can provide historical context for weather modification usage in the US. In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of weather modification usage[4].
Methods
Figure 1: NOAA Data Extraction Pipeline using LLM
Input Data Source
Cloud seeding reports were manually downloaded from the NOAA Central Library’s Weather Modification Project Reports archive [2]. A total of 1,025 PDF files were downloaded, covering all reported weather modification activities in the United States from 2000-2025.
Each PDF contains the initial report on the planned cloud seeding project (Form 17-4), a second tabular report describing interim and final details (Form 17-4A), or a combination of both. Form 17-4 contains project-level summaries, including information on the purpose of the project (e.g., precipitation enhancement or hail suppression), the operational period, and the agents and apparatus used. Form 17-4A provides more granular operational details such as the number of seeding hours, the type and amount of agent released, and meteorological conditions on specific dates.
Beyond these two forms, many reports include additional supporting materials such as operational maps, meteorological analyses, auxilliary tables, cloud seeding decision logs, and narrative summaries. While these materials provide valuable context for understanding cloud seeding activities, they also increase variability across reports, requiring additional preprocessing and file handling.
Preprocessing
A custom Python pipeline was developed to organize and preprocess the PDF documents (Figure 1). First, all downloaded files were stored in a local directory. In many cases, a single cloud seeding project was documented across multiple submissions, indicated by suffixes such as .M (initial report), .I (interim report), and .F (final report). Since these reports refer to the same geographic area and operational period, they were programmatically consolidated to preserve project-level information and ensure completeness during data extraction. After consolidation, there were 836 unique weather modification projects to process and include in our dataset.
The scanned documents included both typed and hand-written text (and sometimes a mix of both in the same document). To convert the merged PDFs into plain text, three approaches were used. For native, text-based documents (57% of documents), text was extracted using the pymupdf Python library [5]. For scanned or hand-written documents (43% of documents), optical character recognition (OCR) was applied using a combination of pytesseract, an open-source wrapper for the Tesseract OCR engine [6] (processed 16% of all documents), and llm-whisperer (processed 27% of all documents), a layout-preserving PDF extraction technology which uses both OCR and text-based extraction [7]. All code for the preprocessing pipeline, including file merging and text extraction, is available in a public GitHub repository [8].
For each document, the text from Form 17-4 was extracted using a three-tie approach: first attempting pymupdf, then pytesseract, and lastly llm-whisperer.
Only Form 17-4 was used because it contained all the key metadata we needed for our dataset, including: purpose, date ranges, description of agent and apparatus used, and the target area. Truncating the input in this way also helped reduce noise and minimize costs. In addition to the extracted text, the original filename was also provided to the model because the filenames often contained useful contextual clues about the report’s location or time frame (e.g. North Dakota Cloud Modification Project - District I_02-1147_06.01.2002-08.31.2002.pdf).
Notably, there appeared to be two distinct naming conventions in the archive. Older files (generally before 2017) tend to use a descriptive format that includes the location and date (e.g., Colorado-SanJuan-2003.pdf), whereas newer files (from around 2018 onward) follow a standardized alphanumeric format such as 2018TXSWMA-1_2018.pdf, which embeds the year, state abbreviation, and project code. These filenames were leveraged during preprocessing and extraction to assist with project disambiguation and metadata recovery.
LLM Integration
After experimenting with various model options, we decided to use OpenAI’s o4-mini model due to its balance of performance and cost [9]. We also experimented with three different prompt templates, and the final prompt was chosen to use chain-of-thought reasoning with example reasoning steps to extract fields aligned with Form 17-4 and return structured text outputs for subsequent parsing. All prompt versions and evaluation scripts are available in the GitHub repository [8].
LLM Prompt
# NOAA Weather Modification Report Extraction Expert
You are an expert data extractor specialized in parsing historical NOAA Weather Modification Activity reports. For each report, utilize the PDF-converted text and filename to extract 9 critical fields.
## Instructions
For each field, carefully analyze all available information using a step-by-step reasoning process. Clearly document your reasoning, resolve conflicting or ambiguous information logically, and then provide your final extracted value using the Final Extracted Fields Format.
## Explanation of Fields
The fields you must extract are detailed below, with guidelines for each:
... truncated for brevity ...
## Example Chain-of-Thought Reasoning
**YEAR OF WEATHER MODIFICATION ACTIVITY:**
- Filename: ‘2018UTNORT-1.pdf
- Filename starts with ‘2018‘, clearly indicating the year.
- No conflicting year mentioned elsewhere.
- Final inference: **2018**
... truncated for brevity ...
[Continue similarly structured reasoning for all fields]
## Final Extracted Fields Format
... truncated for brevity ...
Postprocessing
The LLM response was a list containing the key information we requested to extract:
YEAR: [extracted value]
SEASON: [extracted value]
STATE: [extracted value]
AGENT: [extracted value]
APPARATUS: [extracted value]
PURPOSE: [extracted value]
TARGET AREA: [extracted value]
START DATE: [extracted value]
END DATE: [extracted value]
Then, in order to neatly store this information into CSV format, we mapped key values to more concise column names, stripped any blank space, and converted to lower case. The resulting data object now looked like this:
parsed_gpt_response = {’start_date’: ’10/15/2007’, end_date’: ’12/24/2007’, season’: winter’, target_area’: alta and snowbird ski resorts’, year’: ’2007’, state’: utah’, type_of_agent’: silver iodide’, type_of_apparatus’: ground’, purpose’: augment snowpack’, filename’: Alta - Snowbird, Utah_07-1384_10.15.2007-12.24.2007.pdf’}
With the formatted data object, we could easily save the data to CSV with each object as a separate row in the CSV, and each key-value pair corresponding to the columns.
Data Records
The final cloud seeding dataset is provided as a single CSV file titled cloud_seeding_us_2000_2025.csv, using one row per cloud seeding project and one column per extracted field. It is publicly available on Zenodo at the following DOI: https://doi.org/10.5281/zenodo.14925812. It contains structured information extracted from 836 cloud seeding project reports in the United States between 2000 and 2025.
Figure 2: Dataset Screenshot
The following fields are present in the dataset:
  • filename: The original NOAA report filename from which the data was extracted (e.g., 2021TXSWMA_STWMA_2021.pdf).
  • year: The calendar year in which the cloud seeding activity occurred (e.g., 2015).
  • season: The meteorological or operational season during which the activity took place (e.g., winter, summer).
  • state: The U.S. state where the cloud seeding operation was conducted (e.g., california)
  • agent: The material used to seed clouds (e.g., silver iodide, dry ice).
  • apparatus: The method or equipment used to deliver the seeding agent (e.g., ground, airborne).
  • purpose: The stated objective of the cloud seeding activity (e.g., augment snowpack, increase rain).
  • target_area: The geographic area targeted for cloud seeding (e.g., lake tahoe and truckee river basin).
  • start_date: The reported start date of the cloud seeding activity (e.g., 11/15/2014).
  • end_date: The reported end date of the cloud seeding activity (e.g., 4/15/2015)
Data Visualizations
Between 2000 and 2025, most U.S. cloud seeding activity occurred in Utah, Colorado, and eastern California, primarily aimed at enhancing snowpack through ground-based silver iodide dispersion. This pattern is consistent with the nine western states known to have operational cloud seeding programs: California, Colorado, Idaho, Nevada, New Mexico, North Dakota, Texas, Utah, and Wyoming [4]. In Texas, by contrast, efforts focus on increasing rainfall and suppressing hail using a combination of airborne silver iodide dispersion and ground-based shockwave generation. Overall, cloud seeding activity declined between 2002 and 2023, but showed an uptick in 2024, with 2025 also on track for increased activity.
Figure 3: U.S. Cloud Seeding Activity by State (2000-2025)
Figure 4: Cloud Seeding Activity over Time (2000-2025)
Figure 5: Purpose of Cloud Seeding Activity
Technical Validation
A key challenge in using LLMs for data extraction is their generative nature. Although LLMs are capable of producing fluent and structured text, they can generate information that is not present in the source material. In this context, such hallucinated content may include incorrect years, locations, or project descriptions, which could compromise the reliability of downstream analyses involving historical weather modification activities in the United States.
Additional difficulties arise from the quality of the input documents. Many PDFs contain handwritten annotations, low-resolution scans, or incomplete text. For example, some entries refer the reader to information from a previous year or external attachments rather than providing complete answers in the designated fields. These issues reduce the clarity and completeness of the source data, which in turn affects extraction accuracy.
A common source of error involved missing form data. In several cases, including reports from the North Dakota Cloud Modification Project and the Western Kansas Weather Modification Program, submitters left fields blank and instead referred to separate documents such as an operational plan or an attached sheet. These references could not be processed directly, limiting the ability of the model to extract a full set of structured values.
To assess and improve the accuracy of the extracted dataset, we conducted a structured technical validation. This involved comparing the model’s outputs against a set of manually annotated ground truth records, allowing us to quantify field-level accuracy and identify common failure cases across the dataset.
Accuracy Evaluations
To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM. The review was manually conducted independently by two human experts. A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level. For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-ground-truth.py.
Table 1: Field-Level Accuracy (n=200)
Field Accuracy (%)
Year 100.00%
Season 87.94%
State 100.00%
Agent 89.95%
Apparatus 92.96%
Purpose 97.49%
Target Area not annotated
Start Date not annotated
End Date not annotated
Overall Average 94.72%
Model Comparisons
We experimented with various OpenAI models to improve accuracy, including: gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3. This allowed us to observe how accuracy changed with model complexity and specialization. Ultimately, OpenAI’s ”more affordable reasoning model”, o4-mini, achieved the best balance of performance and cost. When comparing the different model variants, the same prompt was used (Prompt C).
Table 2: Model Comparison (n=50) using Prompt C
LLM Accuracy (%)
gpt-4.1 88.33%
gpt-4.1-mini 83.67%
gpt-4o-mini 86.67%
o3 90.33%
o4-mini 88.00%
Prompt Engineering
We also evaluated how different prompts influenced extraction accuracy, holding the model constant with gpt-4.1-mini. We compared three prompt formats: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and chain-of-thought (Prompt C) that encouraged step-by-step reasoning. Prompt C, using chain-of-thought reasoning, substantially outperformed the others, achieving an accuracy of 87.67%. This suggests that explicit reasoning instructions with examples can improve structured data extraction accuracy.
Table 3: Prompt Comparison (n=50) using gpt-4.1-mini
Prompt Accuracy (%)
Prompt A (Detailed) 72.00%
Prompt B (Concise) 57.00%
Prompt C (Chain-of-Thought) 87.67%
Together, these validation efforts demonstrate that LLM-based data extraction can be improved with careful model selection and prompt design. However, while we tested several variants, we did not exhaustively test every permutation of models-and-prompts listed. Further experimentation might reveal additional insights.
Usage Notes
All weather modification reports are self-reported by project operators and may not represent a complete record of weather modification activity (e.g., if there are omissions). For PDF-to-text conversion, note that llm-whisperer is free only for up to 100 PDF page extractions per day. Additionally, note that the OpenAI models referenced (e.g., gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3), incur usage costs depending on model and volume, so users should consult current OpenAI pricing for budgeting larger-scale replication [10].
Code Availability
Python code for extracting the dataset from the NOAA reports can be obtained from the public repository at https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor. Python code for analyzing and visualizing the dataset is available at: https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor.
Acknowledgments
Both authors acknowledge funding from Columbia University’s Data Science Institute. K.D.L. acknowledges funding from the Zegar Family Foundation.

### v2

@@
 Abstract
-We present a structured dataset of reported cloud seeding activities in the U.S.
+Cloud seeding, a weather modification technique aimed at locally enhancing precipitation, has been used in the Western United States since the 1940’s.
+However, datasets to facilitate large-scale analysis of cloud seeding efforts are not currently available.
+To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S.
 from 2000 to 2025, including the year, season, state, seeding agent, apparatus used for deployment, and purpose.
@@
 The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo.
-Our results address the gap in accessible cloud seeding data in the United States, and demonstrate the potential of LLMs to extract structured environmental data from historical documents.
+Our results address the gap in accessible cloud seeding data in the United States, and demonstrate the potential for LLMs to extract structured environmental data from historical documents.
 More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
@@
 These reports are submitted to NOAA via email and compiled on the NOAA Weather Modification website [2].
-Currently, NOAA’s publicly available records on their website are available for 2000-2025.
-We expected earlier records to be available and submitted a data request to NOAA for documents covering 1972–1999, but have not received a response to date.
-As a result, the dataset currently covers 2000-2025.
-If older cloud seeding reports from 1972–1999 are recovered and made publicly available, the same pipeline can be applied with minimal modifications to expand the dataset’s coverage.
 While the NOAA reports are made publicly available online, they are stored individually as scanned PDFs with inconsistent formatting and structure.
@@
 As a result, data on cloud seeding in the U.S., such as locations, dates, purposes, cloud seeding agent, and agent deployment methods, remains largely inaccessible for analysis.
-To solve this data access problem, we created a structured dataset from 836 NOAA cloud seeding reports spanning from 2020 to 2025.
+To solve this data access problem, we created a structured dataset from 836 NOAA cloud seeding reports spanning from 2000 to 2025.
 Our work transforms these scattered, inconsistently formatted documents into a single, structured CSV file that enables the analysis of U.S.
@@
 This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management—many of which exist in PDF form with highly variable formatting.
-By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest.
+By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest[3].
 Beyond its methodological contributions, the resulting dataset has potential value across a wide range of applications.
 Researchers can use it to study long-term patterns in weather modification practices, analyze the use and evolution of different seeding agents and deployment methods, or assess geographic and seasonal trends in operational activity.
-For example, while there have been numerous recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g.
-the 2017 SNOWIE campaign in Western Idaho[3]), there has been little qualitative analysis of reported historical operational cloud seeding activities.
+For example, while there have been several recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g.
+the 2017 Seeded and Natural Orographic Wintertime clouds: the Idaho Experiment (SNOWIE) campaign in Western Idaho[4]), there has been little qualitative analysis of reported historical operational cloud seeding activities.
 With recent renewed interest in both traditional weather modification technology (e.g.
-the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geo-engineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change, this dataset can provide historical context for weather modification usage in the US.
-In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of weather modification usage[4].
+the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geo-engineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change[5], this dataset can provide historical context for weather modification usage in the US.
+In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of large-scale weather modification[3].
 Methods
-Figure 1: NOAA Data Extraction Pipeline using LLM
+Figure 1: NOAA Data Extraction Pipeline
 Input Data Source
@@
 After experimenting with various model options, we decided to use OpenAI’s o4-mini model due to its balance of performance and cost [10].
-We also experimented with three different prompt templates, and the final prompt was chosen to use chain-of-thought reasoning with example reasoning steps to extract fields aligned with Form 17-4 and return structured text outputs for subsequent parsing.
+We also experimented with three different prompt templates, and the final prompt was chosen to use chain-of-thought reasoning with example reasoning steps to extract fields aligned with Form 17-4 and return structured text outputs for subsequent parsing (shown below).
+Chain-of-thought reasoning guides the model to break down complex tasks into intermediate steps, which improves performance on tasks requiring multi-step inference [11].
 All prompt versions and evaluation scripts are available in the GitHub repository [9].
@@
 ## Final Extracted Fields Format
+Present your final extracted results concisely as follows, in lowercase, comma-separating multiple values when applicable.
+Do not include commentary, explanations, or placeholders. Leave fields blank if truly unknowable after exhausting all inference methods.
 ... truncated for brevity ...
@@
 Data Visualizations
-Between 2000 and 2025, most U.S.
-cloud seeding activity occurred in Utah, Colorado, and eastern California, primarily aimed at enhancing snowpack through ground-based silver iodide dispersion.
-This pattern is consistent with the nine western states known to have operational cloud seeding programs: California, Colorado, Idaho, Nevada, New Mexico, North Dakota, Texas, Utah, and Wyoming [4].
-In Texas, by contrast, efforts focus on increasing rainfall and suppressing hail using a combination of airborne silver iodide dispersion and ground-based shockwave generation.
-Overall, cloud seeding activity declined between 2002 and 2023, but showed an uptick in 2024, with 2025 also on track for increased activity.
-Figure 3: U.S. Cloud Seeding Activity by State (2000-2025)
-Figure 4: Cloud Seeding Activity over Time (2000-2025)
-Figure 5: Purpose of Cloud Seeding Activity
+Between 2000 and 2025, cloud seeding activity in the United States was geographically concentrated in a few key states (Figure 3).
+California, Colorado, Utah, and Texas accounted for the majority of operations, with activity highest in western states that rely on snowpack for water supply.
+These records align with states that maintain active cloud seeding programs.
+Over time, the number of weather modification events peaked in the early to mid-2000s, then declined through the 2010s, before rebounding in 2024 and 2025 (Figure 4).
+This timeline shows that while cloud seeding has remained a consistent tool, its usage has fluctuated, likely due to policy changes, drought conditions, and funding cycles.
+The primary stated purpose of cloud seeding events was to increase snowpack, followed by increasing precipitation and rain (Figure 5).
+Snowpack enhancement is particularly important in mountainous regions where runoff from melting snow supplies reservoirs and irrigation systems.
+Finally, the types of agents and apparatus used vary significantly (Figure 6).
+Silver iodide is by far the most common seeding agent, especially in ground-based and airborne operations.
+Other agents such as calcium chloride and carbon dioxide are used less frequently.
+Ground deployment is the most prevalent method, particularly when using silver iodide, although a mix of airborne and dual (ground+airborne) approaches is also present.
+Figure 3: U.S. Cloud Seeding Activity by State (2000–2025). States with active weather modification programs show the highest number of recorded operations.
+Figure 4: Cloud Seeding Activity by State over Time (2000–2025). The number of activities peaked around 2005, declined gradually, and rose again after 2023.
+Figure 5: Purpose of Cloud Seeding Activity (2000–2025). Snowpack enhancement is the leading purpose, followed by increasing precipitation and rain.
+Figure 6: Agent and Apparatus Used for Cloud Seeding (2000–2025). Silver iodide dominates among agents and is deployed most frequently using ground-based apparatus.
 Technical Validation
 A key challenge in using LLMs for data extraction is their generative nature.
-Although LLMs are capable of producing fluent and structured text, they can generate information that is not present in the source material.
+Although LLMs are capable of producing fluent and structured text, they can generate information that is not present in the source material [12].
 In this context, such hallucinated content may include incorrect years, locations, or project descriptions, which could compromise the reliability of downstream analyses involving historical weather modification activities in the United States.
@@
 A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level.
-For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-ground-truth.py.
-Table 1: Field-Level Accuracy (n=200)
+For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-golden.py.
+Table 1: Field-Level Accuracy (n=200) using o4-mini LLM
 Field Accuracy (%)
@@
 End Date not annotated
-Overall Average 94.72%
+Overall Average Accuracy 94.72%
 Model Comparisons
-We experimented with various OpenAI models to improve accuracy, including: gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3.
-This allowed us to observe how accuracy changed with model complexity and specialization.
-Ultimately, OpenAI’s ”more affordable reasoning model”, o4-mini, achieved the best balance of performance and cost.
-When comparing the different model variants, the same prompt was used (Prompt C).
-Table 2: Model Comparison (n=50) using Prompt C
-LLM Accuracy (%)
-gpt-4.1 88.33%
-gpt-4.1-mini 83.67%
-gpt-4o-mini 86.67%
-o3 90.33%
-o4-mini 88.00%
+We evaluated several OpenAI models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3 [10].
+This comparison allowed us to assess how accuracy varied with model complexity and specialization.
+Among these, o4-mini offered the best balance of performance and cost, achieving 95.00% accuracy at approximately \$0.005 per document.
+While o3 achieved the highest accuracy at 96.33%, it was an order of magnitude more expensive, costing around \$0.05 per document.
+Each accuracy value in Table 2 represents the average field-level accuracy across 50 annotated NOAA reports, using a consistent prompt (Prompt C).
+For each report, model outputs were compared to human-labeled values across all fields, and the proportion of correct fields was averaged across the dataset.
+Although large language models are known to generate outputs probabilistically, we observed stable performance across repeated runs of the same model.
+Table 2: Model Comparison (n=50)
+LLM Overall Average Accuracy (%)
+gpt-4.1 93.33%
+gpt-4.1-mini 93.33%
+gpt-4o-mini 91.67%
+o3 96.33%
+o4-mini 95.00%
 Prompt Engineering
-We also evaluated how different prompts influenced extraction accuracy, holding the model constant with gpt-4.1-mini.
-We compared three prompt formats: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and chain-of-thought (Prompt C) that encouraged step-by-step reasoning.
-Prompt C, using chain-of-thought reasoning, substantially outperformed the others, achieving an accuracy of 87.67%.
-This suggests that explicit reasoning instructions with examples can improve structured data extraction accuracy.
-Table 3: Prompt Comparison (n=50) using gpt-4.1-mini
-Prompt Accuracy (%)
-Prompt A (Detailed) 72.00%
-Prompt B (Concise) 57.00%
-Prompt C (Chain-of-Thought) 87.67%
-Together, these validation efforts demonstrate that LLM-based data extraction can be improved with careful model selection and prompt design.
-However, while we tested several variants, we did not exhaustively test every permutation of models-and-prompts listed.
-Further experimentation might reveal additional insights.
+We also evaluated how different prompts influenced extraction accuracy, holding the model constant (o4-mini).
+Three prompt formats were tested: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and a chain-of-thought format (Prompt C) that encouraged step-by-step reasoning.
+Prompt C achieved the highest accuracy at 95.00%, suggesting that explicit reasoning steps can substantially improve structured data extraction.
+Prompt B performed poorly (44.00%) due to frequent missing values in the outputs, likely caused by its lack of field-level guidance and examples.
+In contrast, Prompt A performed better (93.67%) and benefitted from including worked examples for each field.
+While not explicitly labeled as chain-of-thought, Prompt A was influenced by the same principles, demonstrating that even partial integration of reasoning structure can enhance accuracy.
+Table 3: Prompt Comparison (n=50) using o4-mini LLM
+Prompt Overall Average Accuracy (%)
+Prompt A (Detailed) 93.67%
+Prompt B (Concise) 44.00%
+Prompt C (Chain-of-Thought) 95.00%
+Together, these validation efforts demonstrate that LLM-based data extraction can be significantly improved with thoughtful prompt engineering.
+Chain-of-thought reasoning and example-driven prompts, in particular, are well-suited for scientific data extraction from complex archival documents.
+While our evaluation covered key variants, further experimentation could uncover additional insights.
 Usage Notes
-All weather modification reports are self-reported by project operators and may not represent a complete record of weather modification activity (e.g., if there are omissions).
+All weather modification reports in this dataset are self-reported by project operators and may not represent a complete record of U.S.
+weather modification activity, particularly in cases of unsubmitted or omitted filings.
+As of this writing, NOAA’s publicly available records span the years 2000–2025.
+We anticipated earlier records to be accessible and have submitted a formal data request to NOAA for archived documents covering 1972–1999.
+If these older reports are recovered and released, the current extraction pipeline can be applied with minimal modification to extend the dataset’s temporal coverage.
 For PDF-to-text conversion, note that llm-whisperer is free only for up to 100 PDF page extractions per day.
@@
 Python code for extracting the dataset from the NOAA reports can be obtained from the public repository at https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor.
-Python code for analyzing and visualizing the dataset is available at: https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor.
+Python code for analyzing and visualizing the dataset is available at: https://github.com/jdonohue44/US-Cloud-Seeding-Analysis.
 Acknowledgments

### v3

@@
 Abstract
-Cloud seeding, a weather modification technique aimed at locally enhancing precipitation, has been used in the Western United States since the 1940’s.
-However, datasets to facilitate large-scale analysis of cloud seeding efforts are not currently available.
+Cloud seeding, a weather modification technique used to increase precipitation, has been employed in the western United States since the 1940s.
+However, structured datasets are not currently available to analyze these efforts.
 To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S.
@@
 The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo.
-Our results address the gap in accessible cloud seeding data in the United States, and demonstrate the potential for LLMs to extract structured environmental data from historical documents.
+Our results help fill the gap in structured cloud seeding data and demonstrate the potential for LLMs to extract structured environmental data from historical documents.
 More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
@@
 Data Records
-The final cloud seeding dataset is provided as a single CSV file titled cloud_seeding_us_2000_2025.csv, using one row per cloud seeding project and one column per extracted field.
+The final cloud seeding dataset is provided as a single CSV file titled cloud_seeding_us_2000_2025.csv, using one row per cloud seeding project and one column per extracted field (Figure 2).
 It is publicly available on Zenodo at the following DOI: https://doi.org/10.5281/zenodo.14925812.
@@
 Accuracy Evaluations
-To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM.
+To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM (Table 1).
 The review was manually conducted independently by two human experts.
 A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level.
-For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-golden.py.
+For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-golden.py [9].
 Table 1: Field-Level Accuracy (n=200) using o4-mini LLM
@@
 Model Comparisons
-We evaluated several OpenAI models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3 [10].
+We evaluated several OpenAI models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3 [10] (Table 2).
 This comparison allowed us to assess how accuracy varied with model complexity and specialization.
@@
 Prompt Engineering
-We also evaluated how different prompts influenced extraction accuracy, holding the model constant (o4-mini).
+We also evaluated how different prompts influenced extraction accuracy, holding the model constant with o4-mini (Table 3).
 Three prompt formats were tested: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and a chain-of-thought format (Prompt C) that encouraged step-by-step reasoning.
@@
 Together, these validation efforts demonstrate that LLM-based data extraction can be significantly improved with thoughtful prompt engineering.
-Chain-of-thought reasoning and example-driven prompts, in particular, are well-suited for scientific data extraction from complex archival documents.
-While our evaluation covered key variants, further experimentation could uncover additional insights.
+In particular, chain-of-thought reasoning and example-driven prompts are well-suited for scientific data extraction from complex historical documents.
 Usage Notes

### v4

@@
 To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S.
-from 2000 to 2025, including the year, season, state, seeding agent, apparatus used for deployment, and purpose.
-Using OpenAI’s o4-mini large language model (LLM), combined with multi-stage PDF-to-text conversion and response-parsing code, we processed 836 historical reports from the National Oceanic and Atmospheric Administration (NOAA) to extract the data.
-The resulting dataset achieved 94.72% human-verified accuracy across all fields and is publicly available on Zenodo.
+from 2000–2025, including the project name, year, season, state, operator, seeding agent, apparatus used for deployment, stated purpose, target area, control area, start date, and end date.
+Using OpenAI’s o3 large language model (LLM), combined with our multi-stage PDF-to-text extraction pipeline, we processed 832 historical reports from the National Oceanic and Atmospheric Administration (NOAA).
+The resulting dataset demonstrates 98.38% estimated accuracy, based on manual review of a random sample of 200 records, and is publicly available on Zenodo.
 Our results help fill the gap in structured cloud seeding data and demonstrate the potential for LLMs to extract structured environmental data from historical documents.
@@
 Background & Summary
-Cloud seeding is a weather modification technique used to enhance precipitation, typically in regions affected by drought, water scarcity, or to support snowpack accumulation.
-In the United States, cloud seeding has been practiced since the 1940s but reporting was not standardized until the Weather Modification Reporting Act of 1972 [1].
+Cloud seeding is a weather modification technique used to enhance precipitation, typically in regions experiencing drought, water scarcity, or to support snowpack accumulation.
+In the United States, cloud seeding experiments began in the 1940s [1], and systematic reporting was introduced with the Weather Modification Reporting Act of 1972 [2].
 The Act requires individuals and organizations to notify the U.S.
-Department of Commerce at least 10 days before and after conducting weather modification operations.
-These reports are submitted to NOAA via email and compiled on the NOAA Weather Modification website [2].
+Department of Commerce at least 10 days before and after conducting weather modification activities, using Form 17-4 (Initial Report on Weather Modification Activities) and Form 17-4A (Interim Activity Reports and Final Report).
+These forms are intended to be submitted to NOAA and archived on the NOAA Weather Modification website [3].
+However, only reports from 2000 to 2025 are currently available online.
+Records from earlier years are not accessible through NOAA’s website and have not yet been located elsewhere.
+As a result, our dataset covers only the 2000–2025 period, representing the full set of publicly accessible reports to date.
 While the NOAA reports are made publicly available online, they are stored individually as scanned PDFs with inconsistent formatting and structure.
@@
 As a result, data on cloud seeding in the U.S., such as locations, dates, purposes, cloud seeding agent, and agent deployment methods, remains largely inaccessible for analysis.
-To solve this data access problem, we created a structured dataset from 836 NOAA cloud seeding reports spanning from 2000 to 2025.
+To solve this data access problem, we created a structured dataset from 832 NOAA cloud seeding reports spanning from 2000 to 2025.
 Our work transforms these scattered, inconsistently formatted documents into a single, structured CSV file that enables the analysis of U.S.
 weather modification activities over time.
-The dataset has 94.72% accuracy on all extracted fields when compared with the original PDFs.
+The dataset achieves an estimated 98.38% accuracy across all extracted fields, based on a manual evaluation of 200 randomly sampled records (out of 832), each reviewed by two independent human annotators against the original PDFs.
 This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management—many of which exist in PDF form with highly variable formatting.
@@
 For example, while there have been several recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g.
-the 2017 Seeded and Natural Orographic Wintertime clouds: the Idaho Experiment (SNOWIE) campaign in Western Idaho[4]), there has been little qualitative analysis of reported historical operational cloud seeding activities.
+the 2012 ASCII (AgI Seeding Cloud Impact Investigation) over Sierra Madre range in Wyoming or the 2017 Seeded and Natural Orographic Wintertime clouds: the Idaho Experiment (SNOWIE) campaign in Western Idaho[5, 6]), there has been little qualitative analysis of reported cloud seeding activities.
 With recent renewed interest in both traditional weather modification technology (e.g.
-the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geo-engineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change[5], this dataset can provide historical context for weather modification usage in the US.
-In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of large-scale weather modification[3].
+the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geoengineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change[7], this dataset can provide historical context for weather modification usage in the US.
+In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of weather modification[4].
 Methods
@@
 Input Data Source
-Cloud seeding reports were manually downloaded from the NOAA Central Library’s Weather Modification Project Reports archive [2].
-A total of 1,025 PDF files were downloaded, covering all reported weather modification activities in the United States from 2000-2025.
-Each PDF contains the initial report on the planned cloud seeding project (Form 17-4), a second tabular report describing interim and final details (Form 17-4A), or a combination of both.
-Form 17-4 contains project-level summaries, including information on the purpose of the project (e.g., precipitation enhancement or hail suppression), the operational period, and the agents and apparatus used.
-Form 17-4A provides more granular operational details such as the number of seeding hours, the type and amount of agent released, and meteorological conditions on specific dates.
-Beyond these two forms, many reports include additional supporting materials such as operational maps, meteorological analyses, auxilliary tables, cloud seeding decision logs, and narrative summaries.
-While these materials provide valuable context for understanding cloud seeding activities, they also increase variability across reports, requiring additional preprocessing and file handling.
+Cloud seeding reports were manually downloaded from the NOAA Central Library’s Weather Modification Project Reports Archive [3] (https://library.noaa.gov/weather-climate/weather-modification-project-reports).
+A total of 1,025 PDF files were collected, including Form 17-4, Form 17-4A, and other supplemental documents.
+These files represent all publicly reported weather modification activities submitted to NOAA in the United States from 2000 to 2025.
+The reports are regulatory submissions intended for public access and are in the public domain.
+Their reuse and redistribution are permitted under U.S.
+government policy.
+Each PDF included the initial Form 17-4, a follow-up Form 17-4A, or both.
+Form 17-4 contains project-level summaries, including the purpose of the project, the planned operational period, and the seeding agents and apparatus used.
+Form 17-4A provides more detailed tabular information on seeding activity, such as the number of days seeded per month and the quantities of agent released.
+For this dataset, we focused exclusively on extracting fields from Form 17-4, which contained the majority of the relevant descriptive information.
+While many PDFs also included supporting materials such as maps, meteorological analyses, operational logs, and narrative summaries, these varied widely in structure and content.
+As a result, they were excluded from this version of the dataset due to their inconsistent formatting and the additional preprocessing they would require.
 Preprocessing
-A custom Python pipeline was developed to organize and preprocess the PDF documents (Figure 1).
+A custom Python pipeline was developed to organize and preprocess all Form 17-4 PDF documents (Figure 1).
 First, all downloaded files were stored in a local directory.
@@
 Since these reports refer to the same geographic area and operational period, they were programmatically consolidated to preserve project-level information and ensure completeness during data extraction.
-After consolidation, there were 836 unique weather modification projects to process and include in our dataset.
+After consolidation, there were 832 unique weather modification projects to process and include in our dataset.
 The scanned documents included both typed and hand-written text (and sometimes a mix of both in the same document).
-To convert the merged PDFs into plain text, three approaches were used.
+To convert the PDFs into plain text for further processing, a combination of three technologies was used.
 For native, text-based documents (57% of documents), text was extracted using the pymupdf Python library [8].
-For scanned or hand-written documents (43% of documents), optical character recognition (OCR) was applied using a combination of pytesseract, an open-source wrapper for the Tesseract OCR engine [7] (processed 16% of all documents), and llm-whisperer (processed 27% of all documents), a layout-preserving PDF extraction technology which uses both OCR and text-based extraction [8].
+For scanned or hand-written documents (43% of documents), optical character recognition (OCR) was applied using a combination of pytesseract, an open-source wrapper for the Tesseract OCR engine [9] (processed 16% of all documents), and llm-whisperer (processed 27% of all documents), a layout-preserving PDF extraction technology that combines OCR and native-text extraction [10].
 All code for the preprocessing pipeline, including file merging and text extraction, is available in a public GitHub repository [11].
-For each document, the text from Form 17-4 was extracted using a three-tie approach: first attempting pymupdf, then pytesseract, and lastly llm-whisperer.
-Only Form 17-4 was used because it contained all the key metadata we needed for our dataset, including: purpose, date ranges, description of agent and apparatus used, and the target area.
-Truncating the input in this way also helped reduce noise and minimize costs.
-In addition to the extracted text, the original filename was also provided to the model because the filenames often contained useful contextual clues about the report’s location or time frame (e.g.
+In addition to the extracted text, filenames were also provided to the model because the filenames often contained useful contextual clues about the report’s location or time frame (e.g.
 North Dakota Cloud Modification Project - District I_02-1147_06.01.2002-08.31.2002.pdf).
-Notably, there appeared to be two distinct naming conventions in the archive.
+Notably, there appeared to be two distinct file naming conventions in the archive.
 Older files (generally before 2017) tend to use a descriptive format that includes the location and date (e.g., Colorado-SanJuan-2003.pdf), whereas newer files (from around 2018 onward) follow a standardized alphanumeric format such as 2018TXSWMA-1_2018.pdf, which embeds the year, state abbreviation, and project code.
-These filenames were leveraged during preprocessing and extraction to assist with project disambiguation and metadata recovery.
+These filenames were leveraged during extraction to assist with metadata recovery.
 LLM Integration
-After experimenting with various model options, we decided to use OpenAI’s o4-mini model due to its balance of performance and cost [10].
-We also experimented with three different prompt templates, and the final prompt was chosen to use chain-of-thought reasoning with example reasoning steps to extract fields aligned with Form 17-4 and return structured text outputs for subsequent parsing (shown below).
-Chain-of-thought reasoning guides the model to break down complex tasks into intermediate steps, which improves performance on tasks requiring multi-step inference [11].
-All prompt versions and evaluation scripts are available in the GitHub repository [9].
+We integrated with OpenAI’s platform of LLMs to analyze the extracted text and synthesize the key metadata fields for the structured dataset.
+After evaluating multiple models, the o3 model demonstrated the best performance and was selected for all extractions [12].
 LLM Prompt
+We experimented with multiple prompt templates using prompt engineering techniques to improve extraction quality.
+The final prompt, shown below with truncation, employs chain-of-thought reasoning by guiding the model through intermediate steps before producing structured outputs.
+This reasoning style improves performance on tasks requiring multi-step inference and disambiguation [13].
+The prompt was used with the o3 model for all extractions in the final dataset.
+All prompt versions and evaluation scripts are available in our GitHub repository [11].
 # NOAA Weather Modification Report Extraction Expert
 You are an expert data extractor specialized in parsing historical NOAA Weather Modification Activity reports.
-For each report, utilize the PDF-converted text and filename to extract 9 critical fields.
+For each report, utilize the PDF-converted text and filename to extract **12** critical fields.
 ## Instructions
 For each field, carefully analyze all available information using a step-by-step reasoning process.
-Clearly document your reasoning, resolve conflicting or ambiguous information logically, and then provide your final extracted value using the Final Extracted Fields Format.
-## Explanation of Fields
-The fields you must extract are detailed below, with guidelines for each:
-... truncated for brevity ...
-## Example Chain-of-Thought Reasoning
-**YEAR OF WEATHER MODIFICATION ACTIVITY:**
-- Filename: ‘2018UTNORT-1.pdf
-- Filename starts with ‘2018‘, clearly indicating the year.
-- No conflicting year mentioned elsewhere.
-- Final inference: **2018**
-... truncated for brevity ...
-[Continue similarly structured reasoning for all fields]
+Reason step-by-step internally, using evidence from the filename and report content to resolve conflicting or ambiguous information.
+Output only the final extracted fields using the format and rules provided below.
+## Fields to Extract
+... [truncated for brevity] ...
+## Example Reasoning (Internal)
+... [truncated for brevity] ...
 ## Final Extracted Fields Format
 Present your final extracted results concisely as follows, in lowercase, comma-separating multiple values when applicable.
-Do not include commentary, explanations, or placeholders. Leave fields blank if truly unknowable after exhausting all inference methods.
-... truncated for brevity ...
-Postprocessing
-The LLM response was a list containing the key information we requested to extract:
+Do not include commentary, explanations, or placeholders.
+Leave field blank if truly unknowable after exhausting all inference methods using the filename and text evidence.
+PROJECT: [extracted value]
 YEAR: [extracted value]
@@
 STATE: [extracted value]
+OPERATOR AFFILIATION: [extracted value]
 AGENT: [extracted value]
@@
 TARGET AREA: [extracted value]
+CONTROL AREA: [extracted value]
 START DATE: [extracted value]
 END DATE: [extracted value]
-Then, in order to neatly store this information into CSV format, we mapped key values to more concise column names, stripped any blank space, and converted to lower case.
-The resulting data object now looked like this:
-parsed_gpt_response = {’start_date’: ’10/15/2007’, end_date’: ’12/24/2007’, season’: winter’, target_area’: alta and snowbird ski resorts’, year’: ’2007’, state’: utah’, type_of_agent’: silver iodide’, type_of_apparatus’: ground’, purpose’: augment snowpack’, filename’: Alta - Snowbird, Utah_07-1384_10.15.2007-12.24.2007.pdf’}
-With the formatted data object, we could easily save the data to CSV with each object as a separate row in the CSV, and each key-value pair corresponding to the columns.
+Postprocessing
+As instructed, the LLM analyzed the extracted PDF text and responded with a list containing the key information we requested.
+In order to neatly store this information into CSV format, we then mapped key values into more concise column names, stripped any blank space, and converted to lowercase.
+Using the formatted data object, each project could be easily written to CSV, with one row per object and columns corresponding to the extracted fields.
+Lastly, to prepare the final dataset, we used clean-dataset.py [11] to remove any duplicates, standardize formatting, and fill blank cells with NaN.
 Data Records
-The final cloud seeding dataset is provided as a single CSV file titled cloud_seeding_us_2000_2025.csv, using one row per cloud seeding project and one column per extracted field (Figure 2).
-It is publicly available on Zenodo at the following DOI: https://doi.org/10.5281/zenodo.14925812.
-It contains structured information extracted from 836 cloud seeding project reports in the United States between 2000 and 2025.
-Figure 2: Dataset Screenshot
+The resulting dataset is publicly available on Zenodo [14] as a single CSV file.
+It contains structured information from 832 unique cloud seeding project reports across the United States from 2000 to 2025, with one row per project and one column per extracted field (Figure 2).
+Figure 2: Dataset Screenshot (cloud_seeding_us_2000_2025.csv)
 The following fields are present in the dataset:
-  • filename: The original NOAA report filename from which the data was extracted (e.g., 2021TXSWMA_STWMA_2021.pdf).
+  • filename: The original name of the file from which the data was extracted.
+  • project: The official name or designation of the cloud seeding project or activity.
   • year: The calendar year in which the cloud seeding activity occurred (e.g., 2015).
   • season: The meteorological or operational season during which the activity took place (e.g., winter, summer).
-  • state: The U.S. state where the cloud seeding operation was conducted (e.g., California)
+  • state: The U.S. state where the cloud seeding operation was conducted (e.g., California).
+  • operator_affiliation: The organization or agency responsible for conducting or overseeing the seeding operation.
   • agent: The material used to seed clouds (e.g., silver iodide, dry ice).
-  • apparatus: The method or equipment used to deliver the seeding agent (e.g., ground, airborne).
-  • purpose: The stated objective of the cloud seeding activity (e.g., augment snowpack, increase rain).
-  • target_area: The geographic area targeted for cloud seeding (e.g., Lake Tahoe and Truckee River Basin).
-  • start_date: The reported start date of the cloud seeding activity (e.g., 11/15/2014).
-  • end_date: The reported end date of the cloud seeding activity (e.g., 4/15/2015)
+  • apparatus: The method or equipment used to deliver the seeding agent (e.g., ground-based generators, aircraft).
+  • purpose: The stated objective of the cloud seeding activity (e.g., augment snowpack, increase rainfall, suppress hail).
+  • target_area: The geographical area where the cloud seeding was intended to take effect.
+  • control_area: A designated area used as a comparison region, typically not subject to cloud seeding.
+  • start_date: The first reported date of seeding activity for the project.
+  • end_date: The final reported date of seeding activity for the project.
 Data Visualizations
-Between 2000 and 2025, cloud seeding activity in the United States was geographically concentrated in a few key states (Figure 3).
-California, Colorado, Utah, and Texas accounted for the majority of operations, with activity highest in western states that rely on snowpack for water supply.
-These records align with states that maintain active cloud seeding programs.
-Over time, the number of weather modification events peaked in the early to mid-2000s, then declined through the 2010s, before rebounding in 2024 and 2025 (Figure 4).
-This timeline shows that while cloud seeding has remained a consistent tool, its usage has fluctuated, likely due to policy changes, drought conditions, and funding cycles.
+Between 2000 and 2025, cloud seeding activity in the United States was geographically concentrated in western states that rely on snowpack for water supply.
+California, Colorado, and Utah accounted for the majority of seeding activities, with summertime rain enhancement in Texas also contributing a substantial number of activities (Figure 3).
+These records align with states that maintain active cloud seeding programs [4].
 The primary stated purpose of cloud seeding events was to increase snowpack, followed by increasing precipitation and rain (Figure 5).
 Snowpack enhancement is particularly important in mountainous regions where runoff from melting snow supplies reservoirs and irrigation systems.
-Finally, the types of agents and apparatus used vary significantly (Figure 6).
-Silver iodide is by far the most common seeding agent, especially in ground-based and airborne operations.
+Silver iodide is by far the most common seeding agent, especially in ground-based and airborne operations (Figure 6).
 Other agents such as calcium chloride and carbon dioxide are used less frequently.
-Ground deployment is the most prevalent method, particularly when using silver iodide, although a mix of airborne and dual (ground+airborne) approaches is also present.
-Figure 3: U.S. Cloud Seeding Activity by State (2000–2025). States with active weather modification programs show the highest number of recorded operations.
+Ground deployment is the most prevalent deployment method, particularly when using silver iodide, although a mix of ground and airborne (ground, airborne) approaches is also present.
+Over this 25 year period, the number of weather modification events peaked in the early to mid-2000s, then declined through the 2010s, before rebounding in 2024 and 2025 (Figure 4).
+Figure 3: U.S.
+Cloud Seeding Activity by State (2000–2025).
+States with active weather modification programs show the highest number of recorded operations.
+Specific locations (shown as salmon markers) are from geocoding target_area and state fields with the GoogleMaps API and have not been verified for accuracy.
 Figure 4: Cloud Seeding Activity by State over Time (2000–2025). The number of activities peaked around 2005, declined gradually, and rose again after 2023.
@@
 Technical Validation
-A key challenge in using LLMs for data extraction is their generative nature.
+To address the potential for LLMs to hallucinate content not present in the source documents, and the inconsistent quality of the input PDFs themselves, we measured field-level extraction accuracy by manually reviewing a random sample of 200 records extracted by the LLM and comparing them to ground truth.
 Although LLMs are capable of producing fluent and structured text, they can generate information that is not present in the source material [15].
-In this context, such hallucinated content may include incorrect years, locations, or project descriptions, which could compromise the reliability of downstream analyses involving historical weather modification activities in the United States.
-Additional difficulties arise from the quality of the input documents.
-Many PDFs contain handwritten annotations, low-resolution scans, or incomplete text.
-For example, some entries refer the reader to information from a previous year or external attachments rather than providing complete answers in the designated fields.
-These issues reduce the clarity and completeness of the source data, which in turn affects extraction accuracy.
-A common source of error involved missing form data.
-In several cases, including reports from the North Dakota Cloud Modification Project and the Western Kansas Weather Modification Program, submitters left fields blank and instead referred to separate documents such as an operational plan or an attached sheet.
+In the context of this paper, such hallucinated content could include incorrect dates, locations, or project descriptions, which would compromise the reliability of downstream analyses involving historical weather modification activities in the United States.
+Additionally, the quality of the input documents presented another challenge.
+Many PDFs contained handwritten annotations, low-resolution scans, or incomplete text referring the reader to information from a previous year or external attachments rather than providing complete answers in the designated fields.
+For example, reports from the North Dakota Cloud Modification Project and the Western Kansas Weather Modification Program left fields blank and instead referred to separate documents such as an operational plan or an attached sheet.
 These references could not be processed directly, limiting the ability of the model to extract a full set of structured values.
-To assess and improve the accuracy of the extracted dataset, we conducted a structured technical validation.
-This involved comparing the model’s outputs against a set of manually annotated ground truth records, allowing us to quantify field-level accuracy and identify common failure cases across the dataset.
 Accuracy Evaluations
-To evaluate field-level extraction accuracy, we manually reviewed a random sample of 200 records (n=200), each consisting of the fields extracted from NOAA weather modification reports by the LLM (Table 1).
-The review was manually conducted independently by two human experts.
-A sample size of 200 was selected to provide a margin of error of ±4% at a 90% confidence level.
-For each record, extracted fields were compared against a ground truth dataset created by the human experts using compare-to-golden.py [9].
-Table 1: Field-Level Accuracy (n=200) using o4-mini LLM
-Field Accuracy (%)
-Year 100.00%
-Season 87.94%
-State 100.00%
-Agent 89.95%
-Apparatus 92.96%
-Purpose 97.49%
-Target Area not annotated
-Start Date not annotated
-End Date not annotated
-Overall Average Accuracy 94.72%
+From the population of N=832 records we randomly sampled n=200 to evaluate field-level extraction accuracy (Table 1).
+This evaluation was conducted by two independent human experts manually creating a golden dataset (ground truth) using the source PDFs, then comparing the extracted fields against the golden dataset using compare-to-golden.py [11].
+Assuming a true population accuracy of 95%, this sample size was chosen to yield a margin of error of $\pm 3$ percentage points at the 95% confidence level.
+Table 1: Field-Level Extraction Accuracy (n=200; model=o3)
+Field Accuracy (%) 95% Confidence Interval
+Project 100.00% [97.00%, 100.00%]
+Year 100.00% [97.00%, 100.00%]
+Season 100.00% [97.00%, 100.00%]
+State 99.50% [96.50%, 100.00%]
+Operator Affiliation 98.50% [95.50%, 100.00%]
+Agent 98.50% [95.50%, 100.00%]
+Apparatus 97.00% [94.00%, 100.00%]
+Purpose 99.50% [96.50%, 100.00%]
+Target Area 97.50% [94.50%, 100.00%]
+Control Area 92.00% [89.00%, 95.00%]
+Start Date 98.50% [95.50%, 100.00%]
+End Date 99.50% [96.50%, 100.00%]
+Overall Average 98.38% [95.38%, 100.00%]
 Model Comparisons
-We evaluated several OpenAI models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3 [10] (Table 2).
-This comparison allowed us to assess how accuracy varied with model complexity and specialization.
-Among these, o4-mini offered the best balance of performance and cost, achieving 95.00% accuracy at approximately \$0.005 per document.
-While o3 achieved the highest accuracy at 96.33%, it was an order of magnitude more expensive, costing around \$0.05 per document.
-Each accuracy value in Table 2 represents the average field-level accuracy across 50 annotated NOAA reports, using a consistent prompt (Prompt C).
-For each report, model outputs were compared to human-labeled values across all fields, and the proportion of correct fields was averaged across the dataset.
-Although large language models are known to generate outputs probabilistically, we observed stable performance across repeated runs of the same model.
-Table 2: Model Comparison (n=50)
+We evaluated several language models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3[12] (Table 2).
+To isolate the effect of model choice, we held the prompt constant using a fixed prompt version (Prompt C [11]) across all evaluations.
+For each trial, model outputs were compared to human-labeled values across all fields, and the proportion of correct fields was averaged across the dataset.
+The o3 reasoning model achieved the highest overall accuracy at 96.33%, but was initially an order of magnitude more expensive than o4-mini costing approximately \$0.05 per document versus \$0.005 for o4-mini, which achieved 95.00% accuracy.
+However, in June 2025, OpenAI reduced the price of o3, bringing its cost down to \$0.01 per document.
+Following this price drop, we adopted o3 to maximize accuracy without increasing cost.
+Table 2: Model Comparison (n=50; prompt C)
 LLM Overall Average Accuracy (%)
@@
 gpt-4o-mini 91.67%
-o3 96.33%
+o3 98.38%
 o4-mini 95.00%
 Prompt Engineering
-We also evaluated how different prompts influenced extraction accuracy, holding the model constant with o4-mini (Table 3).
+To isolate the effect of prompt design, we measured extraction accuracy with a single model, o4-mini (Table 3).
+We chose o4-mini because it offers a favorable balance between cost and performance.
+While the absolute accuracies are specific to o4-mini, we expect the relative ranking of prompts to extend to other models, such as o3.
 Three prompt formats were tested: detailed markdown instructions (Prompt A), concise instructions (Prompt B), and a chain-of-thought format (Prompt C) that encouraged step-by-step reasoning.
@@
 Prompt B performed poorly (44.00%) due to frequent missing values in the outputs, likely caused by its lack of field-level guidance and examples.
-In contrast, Prompt A performed better (93.67%) and benefitted from including worked examples for each field.
+In contrast, Prompt A performed better (93.67%) and benefited from including worked examples for each field.
 While not explicitly labeled as chain-of-thought, Prompt A was influenced by the same principles, demonstrating that even partial integration of reasoning structure can enhance accuracy.
@@
 weather modification activity, particularly in cases of unsubmitted or omitted filings.
+In addition, research field campaigns that piggy-backed on existing cloud seeding operations, such as the ASCII field campaign’s monitoring of the Wyoming Weather Modification Pilot Project[5], may not show up as distinct entries in the data record.
 As of this writing, NOAA’s publicly available records span the years 2000–2025.
-We anticipated earlier records to be accessible and have submitted a formal data request to NOAA for archived documents covering 1972–1999.
-If these older reports are recovered and released, the current extraction pipeline can be applied with minimal modification to extend the dataset’s temporal coverage.
-For PDF-to-text conversion, note that llm-whisperer is free only for up to 100 PDF page extractions per day.
+If older reports (from 1972–1999) are recovered and released, the current extraction pipeline can be applied with minimal modification to extend the dataset’s temporal coverage.
+For PDF-to-text conversion, note that llm-whisperer is free only for up to 100 PDF pages per day.
 Additionally, note that the OpenAI models referenced (e.g., gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3), incur usage costs depending on model and volume, so users should consult current OpenAI pricing for budgeting larger-scale replication [16].
 Code Availability
-Python code for extracting the dataset from the NOAA reports can be obtained from the public repository at https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor.
-Python code for analyzing and visualizing the dataset is available at: https://github.com/jdonohue44/US-Cloud-Seeding-Analysis.
+Python code for extracting the dataset from the NOAA reports can be obtained from the public repository [11].
+Python code for analyzing and visualizing the dataset is also publicly available on GitHub [17].
 Acknowledgments

### v5

@@
-Structured dataset of reported cloud seeding activities in the United States (2000–2025) using a large language model
+Structured dataset of reported cloud seeding activities in the United States (2000–2025) using an LLM
 Abstract
-Cloud seeding, a weather modification technique used to increase precipitation, has been employed in the western United States since the 1940s.
-However, structured datasets are not currently available to analyze these efforts.
+Cloud seeding, a weather modification technique used to increase precipitation, has been practiced in the western United States since the 1940s.
+However, comprehensive datasets are not currently available to analyze these efforts.
 To address this gap, we present a structured dataset of reported cloud seeding activities in the U.S.
 from 2000–2025, including the project name, year, season, state, operator, seeding agent, apparatus used for deployment, stated purpose, target area, control area, start date, and end date.
-Using OpenAI’s o3 large language model (LLM), combined with our multi-stage PDF-to-text extraction pipeline, we processed 832 historical reports from the National Oceanic and Atmospheric Administration (NOAA).
-The resulting dataset demonstrates 98.38% estimated accuracy, based on manual review of a random sample of 200 records, and is publicly available on Zenodo.
-Our results help fill the gap in structured cloud seeding data and demonstrate the potential for LLMs to extract structured environmental data from historical documents.
+Combining our multi-stage PDF-to-text extraction pipeline with OpenAI’s o3 large language model (LLM), we processed 832 historical reports from the National Oceanic and Atmospheric Administration (NOAA).
+The resulting dataset demonstrates 98.38% estimated accuracy, based on manual review of 200 randomly sampled records, and is publicly available on Zenodo.
+This dataset addresses the gap in cloud seeding data and demonstrates the potential for LLMs to extract structured information from historical environmental documents.
 More broadly, this work provides a scalable framework for unlocking historical data from scanned documents across scientific domains.
@@
 weather modification activities over time.
-The dataset achieves an estimated 98.38% accuracy across all extracted fields, based on a manual evaluation of 200 randomly sampled records (out of 832), each reviewed by two independent human annotators against the original PDFs.
-This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management—many of which exist in PDF form with highly variable formatting.
+The dataset achieves an estimated 98.38% accuracy across all extracted fields, based on a manual evaluation of 200 randomly sampled records (out of the total population of 832), each reviewed by two independent human annotators against the original PDFs.
+This data extraction method can be extended to similar government-mandated environmental reporting systems, such as those related to water usage, air quality monitoring, or land management, many of which exist in PDF form with highly variable formatting.
 By transforming previously inaccessible records into structured data, this work lays the groundwork for evidence-based research and policy evaluation in an area of growing environmental interest[4].
@@
 Researchers can use it to study long-term patterns in weather modification practices, analyze the use and evolution of different seeding agents and deployment methods, or assess geographic and seasonal trends in operational activity.
-For example, while there have been several recent research field campaigns that have focused on detailed process level understanding and evaluation of cloud seeding (e.g.
-the 2012 ASCII (AgI Seeding Cloud Impact Investigation) over Sierra Madre range in Wyoming or the 2017 Seeded and Natural Orographic Wintertime clouds: the Idaho Experiment (SNOWIE) campaign in Western Idaho[5, 6]), there has been little qualitative analysis of reported cloud seeding activities.
+For example, while there have been several recent research campaigns focused on process-level understanding and evaluation of cloud seeding, such as the the 2012 ASCII (AgI Seeding Cloud Impact Investigation) over the Sierra Madre range in Wyoming, or the 2017 Seeded and Natural Orographic Wintertime clouds: the Idaho Experiment (SNOWIE) campaign in Western Idaho[5, 6], there has been little quantitative analysis of reported cloud seeding activities.
 With recent renewed interest in both traditional weather modification technology (e.g.
-the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as in geoengineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change[7], this dataset can provide historical context for weather modification usage in the US.
+the UAE’s Rain Enhancement Program and several recent commercial start-ups), as well as geoengineering strategies focused on Solar Radiation Management such as Marine Cloud Brightening that use similar intervention strategies to mitigate climate change[7], this dataset can provide historical context for weather modification usage in the United States.
 In addition, this dataset can provide insights into current limitations in reporting requirements that impede large-scale evaluation of the environmental and meteorological implications of weather modification[4].
 Methods
-Figure 1: NOAA Data Extraction Pipeline
+Figure 1: NOAA Form 17-4 Dataset Extraction Pipeline
 Input Data Source
-Cloud seeding reports were manually downloaded from the NOAA Central Library’s Weather Modification Project Reports Archive [3] (https://library.noaa.gov/weather-climate/weather-modification-project-reports).
+Cloud seeding reports were manually downloaded from the NOAA Central Library’s Weather Modification Project Reports Archive (https://library.noaa.gov/weather-climate/weather-modification-project-reports) [3].
 A total of 1,025 PDF files were collected, including Form 17-4, Form 17-4A, and other supplemental documents.
-These files represent all publicly reported weather modification activities submitted to NOAA in the United States from 2000 to 2025.
-The reports are regulatory submissions intended for public access and are in the public domain.
-Their reuse and redistribution are permitted under U.S.
+These 1,025 files represent all publicly reported weather modification activities submitted to NOAA.
+Since the reports are regulatory submissions intended for public access, they are in the public domain and their reuse and redistribution are permitted under U.S.
 government policy.
-Each PDF included the initial Form 17-4, a follow-up Form 17-4A, or both.
-Form 17-4 contains project-level summaries, including the purpose of the project, the planned operational period, and the seeding agents and apparatus used.
-Form 17-4A provides more detailed tabular information on seeding activity, such as the number of days seeded per month and the quantities of agent released.
-For this dataset, we focused exclusively on extracting fields from Form 17-4, which contained the majority of the relevant descriptive information.
-While many PDFs also included supporting materials such as maps, meteorological analyses, operational logs, and narrative summaries, these varied widely in structure and content.
-As a result, they were excluded from this version of the dataset due to their inconsistent formatting and the additional preprocessing they would require.
+Each PDF included Form 17-4, Form 17-4A, or both, along with supplemental pages.
+Form 17-4 contains the main project summary, including the project purpose, operational period, operator, seeding agents, apparatus for deployment, and target/control areas.
+Form 17-4A provides tabular data on seeding activity, including days seeded per month and agent quantities released.
+For our dataset, we focused exclusively on extracting fields from Form 17-4 which contained the majority of relevant project descriptive information.
+While many PDFs included supporting materials, such as maps, meteorological analyses, operational logs, and narrative summaries, these were excluded due to inconsistent formatting that would require extensive preprocessing with minimal dataset enhancement.
 Preprocessing
-A custom Python pipeline was developed to organize and preprocess all Form 17-4 PDF documents (Figure 1).
+A custom Python pipeline was developed to preprocess all Form 17-4 PDF documents (Figure 1).
 First, all downloaded files were stored in a local directory.
 In many cases, a single cloud seeding project was documented across multiple submissions, indicated by suffixes such as .M (initial report), .I (interim report), and .F (final report).
-Since these reports refer to the same geographic area and operational period, they were programmatically consolidated to preserve project-level information and ensure completeness during data extraction.
-After consolidation, there were 832 unique weather modification projects to process and include in our dataset.
-The scanned documents included both typed and hand-written text (and sometimes a mix of both in the same document).
+These reports were consolidated into a single file per project using a combination of programmatic and manual methods to preserve project-level information.
+After consolidation, 832 unique weather modification projects remained for processing and inclusion in the dataset.
+The scanned documents included both typed and hand-written text, and sometimes a mix of both in the same document.
 To convert the PDFs into plain text for further processing, a combination of three technologies was used.
@@
 For scanned or hand-written documents (43% of documents), optical character recognition (OCR) was applied using a combination of pytesseract, an open-source wrapper for the Tesseract OCR engine [9] (processed 16% of all documents), and llm-whisperer (processed 27% of all documents), a layout-preserving PDF extraction technology that combines OCR and native-text extraction [10].
-All code for the preprocessing pipeline, including file merging and text extraction, is available in a public GitHub repository [11].
+All code for the preprocessing pipeline, including file merging, text extraction, and LLM-integration, is publicly available on GitHub (https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor) [11].
 In addition to the extracted text, filenames were also provided to the model because the filenames often contained useful contextual clues about the report’s location or time frame (e.g.
@@
 This reasoning style improves performance on tasks requiring multi-step inference and disambiguation [13].
-The prompt was used with the o3 model for all extractions in the final dataset.
-All prompt versions and evaluation scripts are available in our GitHub repository [11].
+The prompt was used with the o3 reasoning model for all extractions in the final dataset.
+All prompt versions and evaluation scripts are available in our public GitHub repository [11].
 # NOAA Weather Modification Report Extraction Expert
@@
 ## Fields to Extract
-... [truncated for brevity] ...
+[truncated for brevity]
 ## Example Reasoning (Internal)
-... [truncated for brevity] ...
+[truncated for brevity]
 ## Final Extracted Fields Format
@@
 Using the formatted data object, each project could be easily written to CSV, with one row per object and columns corresponding to the extracted fields.
-Lastly, to prepare the final dataset, we used clean-dataset.py [11] to remove any duplicates, standardize formatting, and fill blank cells with NaN.
+Lastly, to prepare the final dataset, we used clean-dataset.py in [11] to remove any duplicates, standardize formatting, and fill blank cells with NaN.
 Data Records
-The resulting dataset is publicly available on Zenodo [14] as a single CSV file.
+The dataset is available on Zenodo as a single CSV file (https://doi.org/10.5281/zenodo.14925811) [14].
 It contains structured information from 832 unique cloud seeding project reports across the United States from 2000 to 2025, with one row per project and one column per extracted field (Figure 2).
-Figure 2: Dataset Screenshot (cloud_seeding_us_2000_2025.csv)
+Figure 2: Screenshot of the CSV Dataset (cloud_seeding_us_2000_2025.csv)
 The following fields are present in the dataset:
@@
 Over this 25 year period, the number of weather modification events peaked in the early to mid-2000s, then declined through the 2010s, before rebounding in 2024 and 2025 (Figure 4).
-Figure 3: U.S.
-Cloud Seeding Activity by State (2000–2025).
+Figure 3: Cloud Seeding Activity by U.S.
+State (2000–2025).
 States with active weather modification programs show the highest number of recorded operations.
-Specific locations (shown as salmon markers) are from geocoding target_area and state fields with the GoogleMaps API and have not been verified for accuracy.
-Figure 4: Cloud Seeding Activity by State over Time (2000–2025). The number of activities peaked around 2005, declined gradually, and rose again after 2023.
-Figure 5: Purpose of Cloud Seeding Activity (2000–2025). Snowpack enhancement is the leading purpose, followed by increasing precipitation and rain.
-Figure 6: Agent and Apparatus Used for Cloud Seeding (2000–2025). Silver iodide dominates among agents and is deployed most frequently using ground-based apparatus.
+Specific locations (shown as salmon markers) are from geocoding the stated target_area field with the GoogleMaps API.
+Figure 4: Cloud Seeding Activity by U.S. State over Time (2000–2025). The number of activities peaked between 2003-2005, declined gradually, and rose again after 2021.
+Figure 5: Stated Purpose of Cloud Seeding Activity (2000–2025). Augmenting snowpack is the leading purpose, followed by increasing precipitation.
+Figure 6: Agent and Apparatus Used for Cloud Seeding Operations (2000–2025). Silver iodide dominates among agents and is deployed most frequently using a ground-based apparatus.
 Technical Validation
 To address the potential for LLMs to hallucinate content not present in the source documents, and the inconsistent quality of the input PDFs themselves, we measured field-level extraction accuracy by manually reviewing a random sample of 200 records extracted by the LLM and comparing them to ground truth.
+This evaluation provides an empirical bound on potential extraction errors in the released dataset, enabling downstream users to incorporate measurement uncertainty into their analyses.
 Although LLMs are capable of producing fluent and structured text, they can generate information that is not present in the source material [15].
@@
 These references could not be processed directly, limiting the ability of the model to extract a full set of structured values.
-Accuracy Evaluations
+Extraction Accuracy
 From the population of N=832 records we randomly sampled n=200 to evaluate field-level extraction accuracy (Table 1).
-This evaluation was conducted by two independent human experts manually creating a golden dataset (ground truth) using the source PDFs, then comparing the extracted fields against the golden dataset using compare-to-golden.py [11].
-Assuming a true population accuracy of 95%, this sample size was chosen to yield a margin of error of $\pm 3$ percentage points at the 95% confidence level.
+This sample size was chosen to yield a margin of error of $\pm 3$ percentage points at the 95% confidence level.
+The evaluation was conducted by two independent human experts manually creating a golden dataset (ground truth) using the source PDFs, then comparing the extracted fields against the golden dataset using compare-to-golden.py [11].
 Table 1: Field-Level Extraction Accuracy (n=200; model=o3)
@@
 Overall Average 98.38% [95.38%, 100.00%]
-Model Comparisons
+LLM Comparisons
 We evaluated several language models to improve extraction accuracy, including gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3[12] (Table 2).
@@
 However, in June 2025, OpenAI reduced the price of o3, bringing its cost down to \$0.01 per document.
-Following this price drop, we adopted o3 to maximize accuracy without increasing cost.
-Table 2: Model Comparison (n=50; prompt C)
-LLM Overall Average Accuracy (%)
-gpt-4.1 93.33%
-gpt-4.1-mini 93.33%
-gpt-4o-mini 91.67%
-o3 98.38%
-o4-mini 95.00%
+Following this price drop, we adopted o3 to maximize accuracy while maintaining low cost.
+Table 2: Extraction Accuracy by LLM (n=50; prompt C)
+LLM Accuracy (%) 95% Confidence Interval
+gpt-4.1 93.33% [86.11%, 97.48%]
+gpt-4.1-mini 93.33% [86.11%, 97.48%]
+gpt-4o-mini 91.67% [84.03%, 96.34%]
+o3 96.33% [88.08%, 98.67%]
+o4-mini 95.00% [88.08%, 98.67%]
 Prompt Engineering
@@
 While not explicitly labeled as chain-of-thought, Prompt A was influenced by the same principles, demonstrating that even partial integration of reasoning structure can enhance accuracy.
-Table 3: Prompt Comparison (n=50) using o4-mini LLM
-Prompt Overall Average Accuracy (%)
-Prompt A (Detailed) 93.67%
-Prompt B (Concise) 44.00%
-Prompt C (Chain-of-Thought) 95.00%
-Together, these validation efforts demonstrate that LLM-based data extraction can be significantly improved with thoughtful prompt engineering.
-In particular, chain-of-thought reasoning and example-driven prompts are well-suited for scientific data extraction from complex historical documents.
+Table 3: Accuracy by LLM Prompt (n=50) using o4-mini
+Prompt Accuracy (%) 95% Confidence Interval
+Prompt A 93.67% [86.64%, 97.56%]
+Prompt B 44.00% [30.98%, 57.69%]
+Prompt C 95.00% [88.08%, 98.67%]
+Together, these validation efforts show that LLM-based data extraction accuracy can be significantly improved through deliberate prompt design and model selection.
+In particular, prompts incorporating chain-of-thought reasoning and field-specific examples consistently enhanced extraction accuracy, even in complex historical documents with heterogeneous formatting and quality.
+These findings suggest that careful prompt engineering should be considered a core methodological step when using LLMs to extract scientific data from historical documents.
 Usage Notes
@@
 Additionally, note that the OpenAI models referenced (e.g., gpt-4.1, gpt-4.1-mini, gpt-4o-mini, o4-mini, and o3), incur usage costs depending on model and volume, so users should consult current OpenAI pricing for budgeting larger-scale replication [16].
+Data Availability
+The dataset is publicly available on Zenodo (https://doi.org/10.5281/zenodo.14925811) [14].
+The original NOAA files used as input are publicly available on NOAA (https://library.noaa.gov/weather-climate/weather-modification-project-reports) [3].
 Code Availability
-Python code for extracting the dataset from the NOAA reports can be obtained from the public repository [11].
-Python code for analyzing and visualizing the dataset is also publicly available on GitHub [17].
+Python code for extracting the dataset from the NOAA reports can be obtained from our public GitHub repository (https://github.com/jdonohue44/NOAA-Weather-Modification-Forms-LLM-Extractor) [11].
+Python code for visualizing the dataset is also publicly available on GitHub (https://github.com/jdonohue44/US-Cloud-Seeding-Analysis) [17].
 Acknowledgments